Please help solve sqrt 15x/4=5
Answer by ankor@dixie-net.com(22740)

:
SqRt(15x/4) = 5
:
Square both sides and you have:
15x/4 = 25
:
Get rid of the denominator, mult eq by 4
15x = 4(25)
15x = 100
x = 100/15
x = 20/3
:
:
Check on good calc: SqRt(15*20/3*1/4) = 5